OpenStudy (anonymous):

9. The sum of two numbers is 51. Their difference is 13. What is the smaller of the two numbers? (Points : 1) 18 19 20 21 10. Charlotte has two job offers as a car sales person. The first job will pay her $600 a month plus 2% on the price of her car sales. The second job will pay $1000 a month plus 1% on the price of her sales. How much must she sell each month to make the same income from the two jobs? (Points : 1) $20,000 $30,000 $40,000 $50,000

OpenStudy (anonymous):

alright for your first question: let x = larger number and let y = smaller number, you have to come up with two equations to solve for x and y, can you think of what those might be?

OpenStudy (anonymous):

19 is the smaller number 31 is the larger

OpenStudy (anonymous):

note the key words of 'sum' and 'difference' in the question

OpenStudy (anonymous):

32 oops

OpenStudy (anonymous):

now that we know ChrisV can give you an answer can we see how you might go about doing it?

OpenStudy (anonymous):

so you can actually learn?

OpenStudy (anonymous):

..just saying doing his hw won't help him on an exam.

OpenStudy (anonymous):

a+b = 5, a= 51-b a-b=13, a=13+b 13+b=51-b 2b=38 b=19

OpenStudy (anonymous):

a+b=51 oops

OpenStudy (anonymous):

better?